A light source emits light of wavelength 125 nm while at rest. When the source is moving, the Doppler shifted wavelength of the emitted light is 375 nm. Is the source of light approaching us or receding from us? Also calculate its speed.



1
Expert's answer
2021-03-24T19:39:19-0400

GIVEN:

"\lambda =125 nm" =emitting wavelength by source


"\lambda'=375nm" =apparent wavelength observed by observer


"c=" 3*108 m\s = speed of light


"v_s=" velocity of source


and apparent wavelength is given by equation,

"\boxed{{\lambda' -\lambda\over \lambda}={v\over c}}"


"v=c*2"

"\boxed{v=6*10^8m/s}"




as the wavelength is increasing therefore, frequency will decrease ...

this implies that the source is receding from us or the observer.